# Does learning the right latent variables necessarily improve in-context learning?

### v1

@@ Abstract
Large autoregressive models like Transformers can solve tasks through in-context learning (ICL) without learning new weights, suggesting avenues for efficiently solving new tasks. For many tasks, e.g., linear regression, the data factorizes: examples are independent given a task latent that generates the data, e.g., linear coefficients. While an optimal predictor leverages this factorization by inferring task latents, it is unclear if Transformers implicitly do so or if they instead exploit heuristics and statistical shortcuts enabled by attention layers. Both scenarios have inspired active ongoing work. In this paper, we systematically investigate the effect of explicitly inferring task latents. We minimally modify the Transformer architecture with a bottleneck designed to prevent shortcuts in favor of more structured solutions, and then compare performance against standard Transformers across various ICL tasks. Contrary to intuition and some recent works, we find little discernible difference between the two; biasing towards task-relevant latent variables does not lead to better out-of-distribution performance, in general. Curiously, we find that while the bottleneck effectively learns to extract latent task variables from context, downstream processing struggles to utilize them for robust prediction. Our study highlights the intrinsic limitations of Transformers in achieving structured ICL solutions that generalize, and shows that while inferring the right latents aids interpretability, it is not sufficient to alleviate this problem. Equal Contribution. Equal Supervision. Correspondence to sarthmit@gmail.com, eric.elmoznino@gmail.com and leogagnon@gmail.com [^1] All code related to our experiments can be found at https://github.com/EricElmoznino/explicit_implicit_icl
@@ 1 Introduction
Recent achievements of large language models (LLMs) showcase the Transformer architecture’s (Vaswani et al., 2017) capacity to extend beyond predictive modeling and solve novel tasks at inference time (Bubeck et al., 2023). However, several papers (Tang et al., 2023; McCoy et al., 2019, 2023) find that rather than modeling the tasks themselves, Transformers rely on shortcuts that risk generalizing poorly to new datasets and tasks. In this work, we aim to shed light on whether learning the true task parameters over shortcuts is sufficient to aid generalization. In particular, we consider the specific case of in-context learning (ICL): the ability of Transformers to leverage demonstrations within their input sequence to adapt to novel queries, effectively learning from contextual cues. ICL underpins much of the capabilities of modern LLMs, including prompt engineering and chain-of-thought, but is difficult to study in LLMs directly due to the multitude of factors that affect performance and lack of control over the training data. To disentangle the role that task latents play in ICL from other aspects that influence large, pre-trained autoregressive models, we therefore introduce controlled experimental settings that involve training Transformers from scratch on tasks that are sufficiently complex, but where the relevant latent variables are well-understood.
Many tasks naturally admit a parametric approach to ICL that breaks the prediction mechanism into two parts: 1) inferring the task-dependent latent variables from the context, and then 2) using them to make predictions on novel queries. For instance, in a linear regression task, a model could first try to infer the underlying weight vector used to generate the data, and subsequently use these inferred weights to make predictions for query points. Another viable methodology, closer in spirit to non-parametric approaches, is to directly compare the query point with the context through a kernel-based mechanism – weighting predictions based on some learned measure of distance between the different points. The two paradigms, parametric vs. non-parametric, reflect tradeoffs between the potential for better generalization or increased flexibility  (Russell & Norvig, 2010).
@@
Recent studies show that Transformers are indeed able to solve ICL tasks through the parametric mechanism in some cases (Hendel et al., 2023; Todd et al., 2024). However, in most cases, there is evidence that Transformers instead rely on non-parametric mechanisms where the prediction is made by directly comparing the query point to exemplars in the context (Wang et al., 2023; Han et al., 2023), a mechanism very much related to induction heads (Olsson et al., 2022). This could be explained by the fact that the functional form of attention operations is almost identical to that of kernel regression (Tsai et al., 2019), making such solutions more natural for Transformers to express (Zhou et al., 2023). These solutions are considered statistical shortcuts since they might not be able to generalize to OOD contexts and queries – for e.g., learning the actual linear predictor for linear regression can generalize to any distribution over training and test points, but nearest-neighbour based interpolation might not.
@@
In this paper, we thereby hypothesize that a limiting factor of ICL in Transformers could be their tendency to prefer non-parametric shortcuts over more structured inference. To test this hypothesis, we perform a thorough analysis through tasks for which latent mechanisms are well understood, and systematically analyse the impact of latent task representation on generalization. We minimally modify the Transformer architecture to prevent such non-parametric shortcuts and compare the OOD performance of the resulting model to that of a traditional Transformer on a large array of synthetic ICL tasks. We call this altered architecture an explicit model in virtue of its inductive bias of explicitly extracting structured latent variables to solve the tasks, and we call the traditional Transformer architecture an implicit model. Specifically, the explicit model prevents the query from directly attending to demonstrations in the context by introducing a dimensionality bottleneck between the processing of the context and the query (see Fig. 1), similar to a conditional neural process (Garnelo et al., 2018a).
Figure 1: We compare the benefits of the implicit (left) and the explicit (right) model. Explicit models disentangle context aggregation and prediction into two separate functions, and have an inductive bias for inferring generative latent variables in order to solve the task. Implicit models are more expressive, but can learn non-parametric shortcut solutions that bypass latent variable inference.
@@
We find that the explicit model does not outperform the implicit one on OOD data, challenging the aforementioned hypothesis that avoiding non-parametric solutions would enhance generalization. Our investigation into this lack of improvement reveals that the issue often lies in the explicit model’s prediction function, which is tasked with leveraging the inferred latent variables for downstream predictions on the query. Our controlled experiments and analysis on the interpretable nature of the bottleneck revealed strong evidence that while the explicit model often extracts relevant task latents, these are not properly utilized by the prediction function.
@@
While on one hand, our research demonstrates that using a simple bottleneck in a Transformer can improve interpretability and explicitly extract task-relevant latent variables, it also suggests that the limitations of Transformers in learning more structured and generalizable ICL solutions are not solely due to non-parametric shortcuts that skirt latent variable inference, but due to more fundamental architectural limitations.
@@
We look at ICL in the context of algorithmic problems where the task is to predict the target $y_{*}$ from a query point ${\bm{x}}_{*}$ when provided with some context examples ${\mathcal{D}}=\{({\bm{x}}_{i},y_{i})\}_{i=1}^{n}$, sharing a common underlying structure defined by the task latent ${\bm{z}}$ and a functional form $g$. The goal of ICL is to learn a function that can utilize the context set ${\mathcal{D}}$ to provide predictions for new query points ${\bm{x}}_{*}$. To achieve this, the model is trained on different draws of context sets (${\mathcal{D}}_{1},{\mathcal{D}}_{2},...$) which share the same underlying functional mapping $g:{\bm{x}},{\bm{z}}\to y$ but different realizations of the latent ${\bm{z}}$, for example $g({\bm{x}},{\bm{z}})={\bm{z}}^{T}{\bm{x}}$ could be a linear regression system shared across different contexts ${\mathcal{D}}_{1},{\mathcal{D}}_{2},...$, but the underlying latents could be different, i.e. ${\mathcal{D}}_{1}$ is generated from ${\bm{z}}_{1}$ while ${\mathcal{D}}_{2}$ from ${\bm{z}}_{2}$, similar to Von Oswald et al. (2023). We emphasize that in this setup, we are not training models to do next-token prediction as is done in language modeling; instead, given a fixed context ${\mathcal{D}}$ that includes $n$ samples, we are attempting to make a prediction on a single novel query ${\bm{x}}_{*}$. We therefore do not use a causal Transformer, and we allow all tokens to attend to each other.
Often, ICL solutions are learned via maximum likelihood, i.e. $\arg\max_{\varphi}\mathbb{E}_{{\mathcal{D}},{\bm{x}}_{*},y_{*}}\left[\log p_{\varphi}(y_{*}|{\bm{x}}_{*},{\mathcal{D}})\right]$, where $p_{\varphi}$ represents the Transformer model and ${\mathcal{D}}$ is sampled from the parametric family defined through $g$. Thus, the transformer model $p_{\varphi}$ must not only learn the form of the prediction function $g$, but also how to efficiently aggregate information from the context ${\mathcal{D}}$ to infer ${\bm{z}}$ for downstream predictions on ${\bm{x}}_{*}$. Thus, this general framework can be naturally decomposed into two distinct parts.
@@
Predictive Modeling. This component refers to the process of estimating the predictive function that leverages context ${\mathcal{D}}$ to infer $y_{*}$ from a query $x_{*}$. In the above example, it refers to learning the functional mapping $g$ once ${\bm{z}}$ has been extracted from context aggregation.
As discussed, Transformers do not have a clear incentive to make this explicit separation of context aggregation and predictive modeling. Instead, given context ${\mathcal{D}}$, they implicitly and jointly model both the function $g$ along with ${\mathcal{D}}$-dependent latent variable ${\bm{z}}$ inference to directly provide predictions for the query point ${\bm{x}}_{*}$, in contrast to separately estimating $g$ and an explicitly factorized ${\bm{z}}$. Thus, in order to enforce explicit representation of ${\bm{z}}$, we propose the simplest architectural modification where the query ${\bm{x}}_{*}$ cannot directly attend to the context, and the latent task representation is forced to summarize the context efficiently. Formally, we compare the following two models, which are illustrated in Fig. 1.
Implicit Model. This refers to the traditional in-context learning computation performed by Transformer models. In this setup, given the set of observations ${\mathcal{D}}$ (context) and a query point ${\bm{x}}_{*}$, the prediction $y_{*}$ is modeled directly as $p_{\varphi}(y_{*}|{\bm{x}}_{*},{\mathcal{D}})$, where $p_{\varphi}$ is defined using a standard Transformer with parameters $\varphi$ and is tasked with modeling both context aggregation and predictive modeling.
@@
Implicit vs. Explicit. Assuming Transformers do in fact favour shortcut-based solutions, we first hypothesize when each setup should perform better given different task characteristics. If the data is generated with a linear model (i.e. $y={\bm{z}}^{T}{\bm{x}}$), the right predictor can be precisely described using the weight vector ${\bm{z}}$, making the explicit model better suited. In contrast, if the data is generated with a Gaussian Process (GP), the implicit model should be superior since by construction query prediction computes similarities with all points in the context. In this case, the task latents of GP-based data with RBF kernel is infinite dimensional (i.e. a point in function space), which cannot be captured in the finite-dimensional bottleneck of the explicit model. In general, we should expect the explicit model to be superior when the underlying true model is parametric and low-dimensional, but in case of a non-parametric or very high dimensional parametric model, the implicit model should be better.
Finally, we note that our aim is not to construct the best possible explicit model architecture – indeed, more sophisticated ones already exist based on amortized Bayesian inference (Garnelo et al., 2018b; Mittal et al., 2023). Instead, we are interested in investigating potential inductive biases for ICL by minimally modifying the standard Transformer architecture to remove certain shortcuts from the space of possible solutions. We leave the design of more performant architectures for future work.
@@ 3 Experiments
Our goal is to use simple tasks that capture the key elements of ICL applications to tease apart the effects of implicit and explicit models on generalization, both in-distribution (ID) and out-of-distribution (OOD).
Task Setup. We conduct experiments across a variety of settings that admit task latents, from synthetic regression and classification to reasoning-based problems. For reasoning tasks that require compositional knowledge, we consider Raven’s Progressive Matrices (Raven’s PM) (John & Raven, 2003), Alchemy (Wang et al., 2021), and Gene Targeting (Norman et al., 2019). A brief description of our tasks is provided below, with a more thorough explanation in §B.
Regression Tasks. We consider different data-generating processes, e.g., linear: prediction ${\bm{z}}^{T}{\bm{x}}$ and latents ${\bm{z}}$, nonlinear (MLP): prediction with a neural network $g({\bm{x}},{\bm{z}})$ and latents as its weights, sinusoidal: prediction as a sum of sinusoids with different frequencies with latensts as their amplitudes, etc.
@@
Raven’s Progressive Matrices. A pattern-completion task used in IQ tests where individual object attributes evolve according to different rules. The task is to complete a sequence of frames such that the underlying rule, which is the latent variable and can be based on colors, shapes, etc., is maintained.
Alchemy. Here, an unknown latent causal graph describes how different stones and potions interact to give rise to new stones. The task is to infer the next stone given a history of transitions.
@@
Training and Evaluation. Tasks based on regression utilize the mean-squared error loss, while those based on classification use the cross entropy loss for training. Model architecture details are provided in §C.1. For ID evaluation, we consider the underlying task latent ${\bm{z}}$, context samples ${\mathcal{D}}$, and queries ${\bm{x}}_{*}$ to be sampled from the same distribution as during training. The challenge in this case is simply to generalize from finite data. For OOD evaluation, we test two different cases depending on the kind of task. For our synthetic regression and classification tasks, the task latent ${\bm{z}}$ and context samples ${\mathcal{D}}$ are sampled from the same distribution as at training time, but the queries ${\bm{x}}_{*}$ are sampled from a Gaussian distribution with higher (3×3$\times$3 ×) standard deviation, thus testing a form of out-of-domain generalization. For our reasoning-based problems, we evaluate on task latents ${\bm{z}}$ that weren’t seen at training. The task latent in each of these reasoning-based problems is composed of parts (e.g., in the Gene Targeting experiment, the latent is a binary vector of targeted genes), which allows us to test a form of compositional generalization (Wiedemer et al., 2023) in which all parts have full marginal support at training time, but novel combinations of those parts are evaluated at test-time.
@@
For all tasks, implicit and explicit Transformer models were trained from scratch over a distribution of tasks latents, and we always control for the number of network parameters to provide a systematic comparison between the implicit and explicit models. To better understand the shortcomings of different models, we also compare with privileged oracles (known decoder – using ground-truth $g$ function, and known latent variable – using an auxiliary loss that includes the ground-truth ${\bm{z}}$).
Figure 2: Comparison of implicit and explicit models both in-distribution (ID) and out-of-distribution (OOD) across a variety of domains: (a) synthetic regression, (b) classification, and (c) compositional generalization tasks. Implicit models are in shown gray, explicit models with Transformer prediction in blue, and with MLP prediction in orange. More details about each task are provided in §B.
@@
Explicit models do not outperform the implicit models. The first evaluation setting that we considered was the ID performance. In this case, we should expect both implicit and explicit models to perform equally well, even if implicit models learn shortcuts rather than ground-truth task latents. This is because those shortcuts are tuned to minimize prediction error within the same data distribution that is being evaluated. Across all our tasks, the results indeed confirmed this prediction. Specifically, during ID evaluation, all models were capable of making highly accurate predictions (Fig. 2). While the performance of the implicit model was generally slightly better than that of the explicit models, the benefits were marginal. Effectively, all models solved the tasks similarly well.
@@
While we expected that ID evaluation would be insufficient to demonstrate potential benefits of the explicit models, we expected to see differences in OOD settings. Both implicit and explicit models are sufficiently expressive to fit the training distribution. However, if an explicit model successfully learns the true task latents that generated the data while an implicit model learns statistical shortcuts that are specialized to the training distribution, we should expect the explicit model to generalize better OOD. As a reminder, for the synthetic regression and classification tasks in Fig. 2 (a, b), OOD evaluation was done by sampling ${\bm{x}}_{*}$ from a normal distribution with wider standard deviation than was used at training (3×)3$\times$)3 × ), effectively evaluating if the models could extrapolate to points further out along their domain despite only being trained within a narrow distribution near the origin. For the compositional tasks in Fig. 2 (c), we instead evaluated OOD performance by only training on certain configurations of the true latent variable ${\bm{z}}$ while evaluating on unseen ones. Importantly, at training time the models were shown data that included every possible value for each component of ${\bm{z}}$, but not ever possible combination of these values were seen, thus evaluating a form of compositional generalization (Wiedemer et al., 2023).

### v2

@@ Abstract
Large autoregressive models like Transformers can solve tasks through in-context learning (ICL) without learning new weights, suggesting avenues for efficiently solving new tasks. For many tasks, e.g., linear regression, the data factorizes: examples are independent given a task latent that generates the data, e.g., linear coefficients. While an optimal predictor leverages this factorization by inferring task latents, it is unclear if Transformers implicitly do so or instead exploit heuristics and statistical shortcuts through attention layers. In this paper, we systematically investigate the effect of explicitly inferring task latents by minimally modifying the Transformer architecture with a bottleneck to prevent shortcuts and incentivize structured solutions. We compare it against standard Transformers across various ICL tasks and find that contrary to intuition and recent works, there is little discernible difference between the two; biasing towards task-relevant latent variables does not lead to better out-of-distribution performance, in general. Curiously, we find that while the bottleneck effectively learns to extract latent task variables from context, downstream processing struggles to utilize them for robust prediction. Our study highlights the intrinsic limitations of Transformers in achieving structured ICL solutions that generalize, and shows that while inferring the right latents aids interpretability, it is not sufficient to alleviate this problem.
Machine Learning, ICML
Figure 1: We compare the benefits of the implicit (left) and the explicit (right) model. Explicit models disentangle context aggregation and prediction into two separate functions with an inductive bias for inferring generative latent variables in order to solve the task. Implicit models are more expressive, but can learn non-parametric shortcut solutions that bypass latent variable inference.
@@ 1 Introduction
Recent advancements in large language models (LLMs, Radford et al., 2019) showcase the Transformer architecture’s (Vaswani et al., 2017) ability to perform novel tasks at inference through in-context learning (ICL, Brown et al., 2020). Indeed, LLMs can learn from instructions and demonstrations provided as input, without requiring gradient-based learning. While ICL plays a key role in many LLM abilities (Lu et al., 2024), such as instruction-following (Wei et al., 2022) and chain-of-thought reasoning (Wei et al., 2023), the factors that influence its generalization – particularly in out-of-distribution (OOD) settings – remain poorly understood. Although ICL can leverage a combination of instructions and demonstrations, our analysis focuses specifically on its ability to model predictions based on a task’s training examples (demonstrations) provided in-context (Lampinen et al., 2024) beyond the modality of language.
A plausible hypothesis behind the success of ICL is that since many tasks are based on some low-dimensional latents (e.g., complex games are described completely through their rules, linear regression through its underlying coefficients), Transformers might generalize to novel queries by inferring the task latents from the context (Hendel et al., 2023; Todd et al., 2024; Yang et al., 2025). This describes a parametric (Bishop, 2006) modeling approach that breaks the prediction mechanism into two parts: 1) inferring the task latents (i.e. parameters) from the context, and then 2) using them to make predictions on novel queries. With the right prediction function, such an approach leads to systematic OOD generalization to new queries.
@@
However, mounting evidence (Wang et al., 2023; Han et al., 2023; Song et al., 2024; Tang et al., 2023b; Bhaskar et al., 2024; Crosbie & Shutova, 2024) suggests that Transformers instead often employ more shallow solutions which rely on direct comparison of the query to demonstrations (reminiscent of induction heads, Olsson et al., 2022). This is closer in spirit to non-parametric approaches (such as nearest neighbors or kernel regression) which are known for their flexibility but poor generalization (Bishop, 2006). Since these solutions do not model the data generative process, they can be described as statistical shortcuts and risk poor performance on OOD context and queries – e.g., learning the actual linear predictor for linear regression can generalize to any distribution over training and test points, but nearest-neighbour based interpolation might not. Interestingly, the functional form of attention operations is almost identical to that of kernel regression (Tsai et al., 2019; Han et al., 2023), making such solutions more natural for Transformers to express (Zhou et al., 2023).
@@
In this paper, we aim to test the hypothesis that biasing Transformers against non-parametric solutions can improve their ICL performance by encouraging parametric modeling. We minimally modify the Transformer architecture to prevent such non-parametric shortcuts and compare the OOD performance of the resulting model to that of a traditional Transformer on a large array of ICL tasks. We call this altered architecture an explicit model because of its inductive bias of explicitly extracting structured latent variables to solve the tasks, and call the traditional Transformer architecture an implicit model. Specifically, the explicit model prevents the query from directly attending to demonstrations in the context by introducing a bottleneck between the processing of the context and the query (see Figure 1), similar to a conditional neural process (Garnelo et al., 2018a). To study the impacts of this inductive bias favoring parametric solutions, we need to establish that explicit models successfully recover task latents. As such, we consider synthetic and real tasks for which the latent mechanisms are well understood, and systematically analyse the impact of task latents on generalization by comparing explicit and implicit models.
@@
We find that the explicit model does not outperform the implicit one on OOD data, challenging the aforementioned hypothesis that avoiding non-parametric solutions enhances generalization. Our investigation into this lack of improvement reveals that the issue often lies in the explicit model’s prediction function which has to leverage the inferred latent variables for downstream predictions on the query. Our controlled experiments and analysis on the interpretable nature of the bottleneck revealed strong evidence that while the explicit model often extracts relevant task latents, these are not properly utilized by the prediction function.
@@
While on one hand, our research demonstrates that using a simple bottleneck in a Transformer can improve interpretability and explicitly extract task-relevant latent variables, it also suggests that the limitations of Transformers in learning more structured and generalizable ICL solutions are not solely due to non-parametric shortcuts that skirt latent variable inference, but due to more fundamental architectural limitations. In sum, our contributions are:
  • Formalizing a framework to test whether parametric ICL solutions generalize better out-of-distribution.
  • Analyzing the benefits, or lack thereof, of inferring the true latents explicitly.
  • Highlighting flaws in the prediction function and downstream latent utilization which hinders generalization.
Figure 2: Comparison of implicit and explicit models in-distribution (ID) and out-of-distribution (OOD) across various domains: (a) synthetic regression, (b) classification, and (c) compositional generalization tasks. Implicit models are in shown gray, explicit models with Transformer prediction in blue, and with MLP prediction in orange. Further details about tasks is provided in Appendix B.
2 Notation
Throughout the paper, we denote datasets with the symbol ${\mathcal{D}}$ which consists of a set of observations with inputs denoted via ${\bm{x}}\in{\mathcal{X}}$ and their corresponding outputs as $y\in{\mathcal{Y}}$. A task is defined by a functional mapping $g:{\mathcal{X}},{\mathcal{Z}}\to{\mathcal{Y}}$ which maps observations ${\bm{x}}$ to labels $y$ through some latents or parameters ${\bm{z}}$, eg. $y={\bm{z}}^{T}{\bm{x}}$ for a linear regression task, or $y\sim{\mathcal{N}}(\cdot;{\bm{z}}^{T}{\bm{x}},\sigma^{2})$ for its stochastic counterpart. To ease readability, we will reserve ${\bm{x}}_{*}\in{\mathcal{X}}$ for the query, i.e. the test time observation we want to generalize to, and $y_{*}\in{\mathcal{Y}}$ its corresponding target. Finally, $\psi$ denotes the parameters of the context aggregation component of explicit model, which inputs the dataset ${\mathcal{D}}$ and infers the corresponding parameters ${\bm{z}}_{\psi}({\mathcal{D}})$, and $\gamma$ the parameters of the prediction model which given a query ${\bm{x}}_{*}$ and parameters ${\bm{z}}\in{\mathcal{Z}}$, provides the prediction. For the implicit model, these operations are subsumed into a single model, with parameters $\varphi$.
@@ 2 Notation
We look at ICL in the context of algorithmic problems where the task is to predict the target $y_{*}$ from a query ${\bm{x}}_{*}$ when provided with some context examples ${\mathcal{D}}=\{({\bm{x}}_{i},y_{i})\}_{i=1}^{n}$, sharing a common underlying structure defined by the task latent ${\bm{z}}$ and a functional form $g$. The goal of ICL is to learn a function that can utilize the context set ${\mathcal{D}}$ to provide predictions for new queries ${\bm{x}}_{*}$. To achieve this, the model is trained on different draws of context sets (${\mathcal{D}}_{1},{\mathcal{D}}_{2},...$) which share the same underlying functional mapping $g:{\bm{x}},{\bm{z}}\to y$ but different realizations of the latent ${\bm{z}}$, for example $g({\bm{x}},{\bm{z}})={\bm{z}}^{T}{\bm{x}}$ could be a linear regression system shared across different contexts ${\mathcal{D}}_{1},{\mathcal{D}}_{2},...$, but the underlying latents could be different, i.e. ${\mathcal{D}}_{1}$ is generated from ${\bm{z}}_{1}$ while ${\mathcal{D}}_{2}$ from ${\bm{z}}_{2}$, similar to Von Oswald et al. (2023). We emphasize that in this setup, we are not training models to do next-token prediction as is done in language modeling; instead, given a fixed context ${\mathcal{D}}$ that includes $n$ samples, we are attempting to make a prediction on a single novel query ${\bm{x}}_{*}$. We therefore do not use a causal Transformer, and we allow all tokens to attend to each other.
Often, ICL solutions are learned via maximum likelihood:
$\arg\max_{\varphi}\;\;\mathbb{E}_{{\mathcal{D}},{\bm{x}}_{*},y_{*}}\left[\log p_{\varphi}(y_{*}|{\bm{x}}_{*},{\mathcal{D}})\right]$ (1)
where $p_{\varphi}$ represents the Transformer model and ${\mathcal{D}}$ is sampled from the parametric family defined through $g$. Thus, the transformer model $p_{\varphi}$ must not only learn the form of the prediction function $g$, but also how to efficiently aggregate information from the context ${\mathcal{D}}$ to infer ${\bm{z}}$ for downstream predictions on ${\bm{x}}_{*}$. Thus, this general framework can be naturally decomposed into two distinct parts.
Figure 3: Performance on a subset of tasks where the true latents ${\bm{z}}$ and prediction function $g$ are known. Implicit models are in shown gray, explicit models with Transformer prediction in blue, models trained with an auxiliary loss to predict the true latents in purple and those using the true prediction function in green. Using the known prediction function leads to significantly better OOD performance.
@@
Predictive Modeling. This component refers to the process of estimating the predictive function that leverages context ${\mathcal{D}}$ to infer $y_{*}$ from a query ${\bm{x}}_{*}$. In the above example, it refers to learning the functional mapping $g$ once ${\bm{z}}$ has been extracted from context aggregation.
As discussed, Transformers do not have a clear incentive to make this explicit separation of context aggregation and predictive modeling. Instead, given context ${\mathcal{D}}$, they implicitly and jointly model both the function $g$ along with ${\mathcal{D}}$-dependent latent variable ${\bm{z}}$ inference to directly provide predictions for the query ${\bm{x}}_{*}$, in contrast to separately estimating $g$ and explicitly factorizing ${\bm{z}}$. Thus, in order to enforce explicit representation of ${\bm{z}}$, we propose a simple architectural modification where the query ${\bm{x}}_{*}$ cannot directly attend to the context, and the latent task representation is forced to summarize the context efficiently. Formally, we compare the following two models, which are illustrated in Figure 1.
Implicit Model. This refers to the traditional in-context learning computation performed by Transformer models. In this setup, given the set of observations ${\mathcal{D}}$ (context) and a query ${\bm{x}}_{*}$, the prediction $y_{*}$ is modeled directly as $p_{\varphi}(y_{*}|{\bm{x}}_{*},{\mathcal{D}})$, where $p_{\varphi}$ is defined using a standard Transformer with parameters $\varphi$ and is tasked with modeling both context aggregation and predictive modeling.
@@
Implicit vs. Explicit. Assuming Transformers do in fact favour shortcut-based solutions, we first hypothesize when each setup should perform better given different task characteristics. If the data is generated with a parametric model with few underlying parameters ${\bm{z}}$ (e.g. a linear model $y={\bm{z}}^{T}{\bm{x}}$), the right predictor can be precisely described using the parameters ${\bm{z}}$, making the explicit model better suited. In contrast, if the data is generated with a Gaussian Process (GP), the implicit model should be superior since by construction query prediction computes similarities with all points in the context. In this case, the task latents of GP-based data with RBF kernel are infinite dimensional (i.e. a point in function space), which cannot be captured in the finite-dimensional bottleneck of the explicit model. In general, we should expect the explicit model to be superior when the underlying true model is parametric and low-dimensional, but in case of a non-parametric or very high dimensional parametric model, the implicit model should be better.
Finally, we note that our aim is not to construct the best possible explicit model architecture – indeed, more sophisticated ones already exist based on amortized Bayesian inference (Garnelo et al., 2018b; Mittal et al., 2023). Instead, we are interested in investigating potential inductive biases for ICL by minimally modifying the standard Transformer architecture to remove certain shortcuts from the space of possible solutions. We leave the design of more performant architectures for future work and refer the readers to Appendix A for a detailed discussion of related work.
Figure 4: Explicit models are interpretable as the bottleneck allows (a) linearly decoding the true latent, and (b) intervening to obtain correct counterfactual predictions. Implicit models are shown in gray, explicit models with Transformer prediction in blue, and with MLP prediction in orange. Models in green use the true prediction function $g$, while models in purple use an additional auxiliary loss based on true latents. To evaluate decoding performance in (a), we linearly decode the true latent directly from concatenated context examples, which yields significantly worse performance than decoding from the bottleneck. Baseline performances in units of the plots are – Linear regression: $0.49$, Nonlinear regression (MLP): $0.94$, Sinusoid regression: $0.33$, Linear classification: $0.86$, Nonlinear classification (MLP): $0.97$, Raven’s PM: $0.5$, and Gene targeting: $0.0$. In (b), the “Relative accuracy” takes the baseline in account (details in C.4).
@@ 4 Experiments
Our goal is to use both synthetic and real tasks that capture the key elements of ICL applications to tease apart the effects of implicit and explicit models on both in-distribution (ID) and out-of-distribution (OOD) generalization.
Task Setup. We conduct experiments across a variety of settings that admit task latents, from synthetic regression and classification to reasoning problems. For reasoning tasks that require compositional knowledge, we consider Raven’s Progressive Matrices (Raven’s PM) (John & Raven, 2003), Alchemy (Wang et al., 2021), Gene Targeting (Norman et al., 2019) and reusable mixture of experts. A brief description of our tasks is provided below, with a more thorough explanation in Appendix B.
Regression Tasks. We consider different data-generating processes, e.g., linear: prediction ${\bm{z}}^{T}{\bm{x}}$ and latents ${\bm{z}}$, nonlinear (MLP): prediction with a neural network $g({\bm{x}},{\bm{z}})$ and latents as its weights, sinusoidal: prediction as a sum of sinusoids with different frequencies and the latents as its amplitudes.
@@
Raven’s Progressive Matrices. A pattern-completion task used in IQ tests where individual object attributes evolve according to different rules. The task is to complete a sequence of frames to satisfy the underlying rule, which is the latent variable and can be based on colors, shapes, etc.
Alchemy. Here, a latent causal graph describes how different stones and potions interact to generate new stones. The task is to infer the next stone given some transitions.
@@
Reusable Modular Mixture of Experts (MoE). In this task, we apply a sequence of operations $g_{l}$ on the input ${\bm{x}}$, where the choice of expert applied at layer $l$ is governed by a categorical latent ${\bm{z}}_{l}$. In particular, we apply five operations sequentially, leading to $y=g_{{\bm{z}}_{5}}\circ g_{{\bm{z}}_{4}}\circ\ldots\circ g_{{\bm{z}}_{1}}({\bm{x}})$. Here $g_{1},g_{2},\ldots$ are shared across tasks and each task is uniquely identified by $({\bm{z}}_{1},\ldots,{\bm{z}}_{5})$ which are the task latents. This represents a reusable mixture of experts system with a hierarchical and compositional decomposition.
@@
Training and Evaluation. Tasks based on regression utilize the mean-squared error loss, while those based on classification use the cross entropy loss for training. Model architecture details are provided in subsection C.1. For ID evaluation, we consider the underlying task latent ${\bm{z}}$, context samples ${\mathcal{D}}$, and queries ${\bm{x}}_{*}$ to be sampled from the same distribution as during training. The challenge in this case is simply to generalize from finite data. For OOD evaluation, we test two different cases depending on the kind of task. For our synthetic regression and classification tasks, the task latent ${\bm{z}}$ and context samples ${\mathcal{D}}$ are sampled from the same distribution as at training time, but the queries ${\bm{x}}_{*}$ are sampled from a Gaussian distribution with higher (3×3$\times$3 ×) standard deviation, thus testing a form of out-of-domain generalization. For our reasoning-based problems, we evaluate on task latents ${\bm{z}}$ that were not seen at training. The task latent in each of these reasoning-based problems is composed of parts (e.g., in the Gene Targeting experiment, the latent is a binary vector of targeted genes), which allows us to test a form of compositional generalization (Wiedemer et al., 2023) in which all parts have full marginal support at training time, but novel combinations of those parts are evaluated at test-time.
@@
For all tasks, implicit and explicit models were trained from scratch over a distribution of tasks latents, with a control for the number of parameters to provide a systematic comparison between the two. To better understand the shortcomings of different models, we also compare with privileged oracles (known decoder – using ground-truth $g$ function, and known latent variable – using an auxiliary loss that includes the ground-truth ${\bm{z}}$).
@@
Explicit models do not outperform the implicit models. The first evaluation setting that we considered was the ID performance. In this case, we should expect both implicit and explicit models to perform equally well, even if implicit models learn shortcuts rather than ground-truth task latents. This is because those shortcuts are tuned to minimize prediction error within the same data distribution that is being evaluated. Across all our tasks, the results indeed confirmed this prediction. Specifically, during ID evaluation, all models were capable of making highly accurate predictions (Figure 2). While the performance of the implicit model was generally slightly better than that of the explicit models, the benefits were marginal. Effectively, all models solved the tasks similarly well.
Figure 5: We conduct experiments on reusable modular MoE task where we train on a subset of combinations of experts, shown on the X-axis. Our results indicate that across different percentages of combinations seen during training, the implicit model consistently outperforms the explicit one at compositional generalization.
@@
While we expected that ID evaluation would be insufficient to demonstrate potential benefits of the explicit models, we expected to see differences in OOD settings. Both implicit and explicit models are sufficiently expressive to fit the training distribution. However, if an explicit model successfully learns the true task latents that generated the data while an implicit model learns statistical shortcuts that are specialized to the training distribution, we should expect the explicit model to generalize better OOD. As a reminder, for the synthetic regression and classification tasks in Figure 2 (a, b), OOD evaluation was done by sampling ${\bm{x}}_{*}$ from a normal distribution with wider standard deviation than was used at training (3×)3$\times$)3 × ), effectively evaluating if the models could extrapolate to points further out along their domain despite only being trained within a narrow distribution near the origin. For the compositional tasks in Figure 2 (c), we instead evaluated OOD performance by only training on certain configurations of the true latent variable ${\bm{z}}$ while evaluating on unseen ones. Importantly, at training time the models were shown data that included every possible value for each component of ${\bm{z}}$, but not every possible combination of these values were seen, thus evaluating a form of compositional generalization (Wiedemer et al., 2023). Similarly, for the reusable modular MoE task, we train the models on a fraction of all possible combinations in the latent $({\bm{z}}_{1},\ldots,{\bm{z}}_{5})$ and evaluate on all combinations.
Figure 6: We analyze (a) Linear regression OOD performance and (b) latent variable linear decoding as a function of model and task parameters. Task performance scales similarly for implicit (gray) and explicit models (blue), while latent variable decoding scales similarly for the explicit model and models trained with the true prediction function $g$ (green).